# How Does a Virtual Agent Decide Where to Look? - Symbolic Cognitive Reasoning for Embodied Head Rotation

### v1

@@ 6.3 Responsiveness to Unseen Distractors
Results are reported in Table 3. Notably, DPS-APC yields an improvement over DPS-DPS, indicating that semi-online validation not only mitigates hallucinations but also enhances responsiveness to scene changes introduced at runtime. Qualitative footage in the supplementary video further illustrates timely reactions to suddenly appearing events, such as an unexpected Santa Claus crossing the field of view.
6.4 Ablation Study
@@ 7 Conclusion
Three limitations suggest directions for future work. First, augmenting perception with spatial audio and haptic streams would allow agents to respond to non-visual cues such as approaching footsteps or vibration alerts. Second, integrating SCORE with a model-predictive navigation controller could unify locomotion and gaze, yielding fully coordinated whole-body behaviour in constrained or collaborative tasks. Third, distilling the LLM into a task-specific lightweight model and caching frequently used predicate patterns in memory may cut inference latency while preserving reasoning depth. Pursuing these extensions could extend SCORE to large-scale social VR, robot telepresence, and other applications that demand believable, responsive head motion.
@@
Figure 6: This example illustrates how SCORE produces human-like head movements in a multi-agent setting. The highlighted agent attends to the intended walking path, aligning closely with the ground-truth behavior. In contrast, Track fixates excessively on static background elements (e.g., storefronts), failing to handle multiple factors (e.g., secure path, understand pedestrian flow) and resulting in implausible head rotations.
@@
Figure 7: Visualization of reasoning and action when Information-seeking,Habit,Safety and Social schema are excluded from the motivational drivers.
Appendix A Implementation details
All experiments were conducted in a Unity environment running on Windows 10. The project was implemented using Unity version 2022.3.55f1. For the Deliberative Perception–Planning Stage (DPS), we employed Gemini 1.5 Pro, a high-performance Large Language Model (LLM) and Vision Language Model (VLM) developed by Google DeepMind, capable of processing both text and images with high accuracy and strong reasoning capabilities. For the Reactive Execution Stage (RES), we employed Gemini 1.5 Flash, a lightweight variant of the model optimized for low-latency inference [43].
Appendix B Single-agent Human dataset details
This supplementary section presents representative samples from our single-agent user study. Each example highlights the object selected by a participant, their corresponding rationale, and a first-person Field-of-View (FoV) image illustrating the interaction. The visualizations are shown in Figure 8,  Figure 9,  Figure 10,  Figure 11, and  Figure 12.
These examples demonstrate that participants not only followed their assigned cognitive goals but also engaged in naturalistic attentional shifts and exploratory behaviors. Each FoV visualization includes a randomly assigned participant ID at the top for identification. The primary object(s) of focus within the scene are highlighted in yellow.
Figure 8: Bus scenario: MDC and APC example. This example demonstrates participant behavior under the cognitive goal of “move to the back." While complying with the instructed task, participants often fixate on contextually salient or socially relevant entities, such as observing the bus driver or scanning for interesting individuals. The yellow highlights indicate the primary objects of attention within the FoV.
Figure 9: Café scenario: MDC and APC example. This example demonstrates participant behavior under the cognitive goal of “find an empty seat and move toward the counter." The yellow highlights indicate the primary objects of attention within the FoV.
Figure 10: Crosswalk scenario: MDC and APC example. This example demonstrates participant behavior under the cognitive goal of “safely cross the street." Participants predominantly focus on safety-relevant cues, such as traffic signals or oncoming vehicles, while intermittently scanning for other engaging elements. The yellow highlights indicate the primary objects of attention within the FoV.
Figure 11: Shopping mall scenario: MDC and APC example. This example demonstrates participant behavior under the cognitive goal of “move toward the escalator." Despite the visual and semantic complexity of the environment, participants effectively maintained goal-directed behavior while attending to notable objects, such as a map for navigation or visually salient entities. The yellow highlights indicate the primary objects of attention within the FoV.
Figure 12: Street scenario: MDC and APC example. This example demonstrates participant behavior under the cognitive goal of “walk safely along the street." Many participants showed heightened attention toward unexpected or socially salient elements, such as an individual lying on the ground. The yellow highlights indicate the primary objects of attention within the FoV.
Appendix C Single-agent head rotation visualization
To intuitively assess how closely our generated head orientations resemble those of humans in a single-agent scenario, we visualize the 3D head rotations over time by projecting them onto a spherical model. Specifically, we extract corresponding data from the human data collected in the user study and compare them with the head trajectory generated by our model, SCORE, as well as by the Track. All time steps of head orientation are plotted as trajectories on a sphere, allowing for clear visual comparison.
Figure 13, Figure 14, Figure 15, Figure 16 and Figure 17 illustrate these visualizations. As shown, the head rotation patterns produced by SCORE demonstrate a strong resemblance to human behavior. This visualization provides compelling qualitative evidence of the alignment between our model’s output and natural human head movements.
Figure 13: Visualization of head rotation examples in the Bus scenario, including both APC and MDC cases. Each plot shows the tracking of the spherical model across all timesteps.
Figure 14: Visualization of head rotation examples in the Café scenario.
Figure 15: Visualization of head rotation examples in the Crossing scenario.
Figure 16: Visualization of head rotation examples in the Mall scenario.
Figure 17: Visualization of head rotation examples in the Street scenario.
Appendix D Multi-agent dataset details
For our experiments, we utilized the UCY crowd dataset [50], a publicly available collection of annotated video sequences widely used for pedestrian trajectory prediction and crowd behavior research. The UCY dataset provides rich annotations of real-world human motion captured in outdoor university settings, using a static overhead camera to simplify perspective distortions and facilitate accurate tracking. Each sequence contains 2D trajectories of pedestrians, manually annotated at a temporal resolution of 2.5 Hz (i.e., one frame every 0.4 seconds).
We specifically employ three sequences from the dataset: Zara01, Zara02, and Students003. Both Zara01 and Zara02 were recorded in a semi-structured pedestrian area adjacent to a shopping complex. These sequences exhibit relatively sparse pedestrian densities, averaging around 5–6 individuals per frame. The scenes include distinct environmental features such as sidewalks, storefronts, doors, walls, and curbs, which naturally guide pedestrian behavior. These sequences capture a wide variety of naturalistic actions, including individuals pausing to look at shop windows, changing direction mid-path, and interacting with environmental obstacles. Due to their clarity and behavioral diversity, these sequences are particularly valuable for extracting examples of nuanced, individualized motion patterns.
In contrast, the Students003 sequence was recorded during a break period outside a university building and represents a dense crowd scenario, with approximately 40 individuals present per frame on average. The setting captures complex inter-agent interactions, such as collision avoidance, local grouping behavior, and flow pattern emergence, making it ideal for studying high-density crowd dynamics. Together, these three sequences offer a complementary set of crowd conditions, ranging from isolated motion to densely packed social navigation, which serve as the foundation for constructing a robust, example-driven crowd simulation model.
Appendix E System prompt Details
Figure 18 illustrates the prompt used in the VLM of the DPS, while Figure 19 shows the corresponding prompt employed in the LLM. Furthermore, Figure 20 presents the prompt designed for the FastVLM model of the RES setting. These prompts were carefully crafted to suit the specific characteristics and tasks of each model, enabling consistent and fair comparisons across different modalities.
Figure 18: The prompt used in the VLM of the DPS.
Figure 19: The prompt used in the LLM of the DPS.
Figure 20: The prompt used in the FastVLM of the RES.

### v2

@@ 6.3. Responsiveness to Unseen Distractors
Results are reported in Table 3. Notably, MDC-APC yields an improvement over MDC-MDC, indicating that semi-online validation not only mitigates hallucinations but also enhances responsiveness to scene changes introduced at runtime. Qualitative footage in the supplementary video further illustrates timely reactions to suddenly appearing events, such as an unexpected Santa Claus crossing the field of view.
6.4. Ablation Study
@@ 7. Conclusion and Future works
We found that SCORE’s rotation decision occasionally diverges from that of humans, reflecting differences not in error but in individual tendencies or persona-like styles. Even under identical conditions, participants displayed substantial variability in where and why they turned their heads, meaning that only probabilistic distributions of plausible behaviors can be inferred rather than a single ground truth trajectory. Similarly, SCORE sometimes produced conclusions that differed from human responses but were still reasonable—often reflecting a more safety-conscious or “diligent” strategy. These results suggest that aligning with human diversity may require modeling variability itself rather than assuming uniformity.
@@
Looking ahead, SCORE holds strong potential for broader applications through further refinement and integration with complementary technologies. First, the current system relies solely on visual inputs, overlooking richer perceptual modalities. Incorporating spatial audio and haptic feedback would enable agents to respond to non-visual cues such as approaching footsteps or vibration alerts. Second, the framework presently operates independently of locomotion control. Coupling SCORE with a model-predictive navigation controller could unify locomotion and gaze, enabling coordinated whole-body behaviors in constrained or collaborative tasks. Third, reliance on large-scale vision–language models introduces inference latency. Distilling the LLM into a task-specific lightweight model and caching frequently used predicate patterns may reduce latency while retaining reasoning depth. Pursuing these extensions could extend SCORE to large-scale social VR, robot telepresence, and other applications that demand believable, responsive head motion.
Acknowledgements.
This work was supported by the Sponsor National Research Foundation of Korea https://www.nrf.re.kr/eng/main under Grant No. Grant #RS-2025-00518643 (34%). It was also partly supported by ICT Creative Consilience Program through the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) under Grant No. IITP-2025-RS-2020-II201819 (33%). It was also supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale under Grant No.RS-2025-02653113 (33%).
@@ Acknowledgements.
Figure 7. This example illustrates how SCORE produces human-like head movements in a multi-agent setting. The highlighted agent attends to the intended walking path, aligning closely with the ground-truth behavior. In contrast, Track fixates excessively on static background elements (e.g., storefronts), failing to handle multiple factors (e.g., secure path, understand pedestrian flow) and resulting in implausible head rotations.
@@
Figure 8. Visualization of reasoning and action when Information-seeking,Habit,Safety and Social schema are excluded from the motivational drivers.